OpenStudy (anonymous):

i need someone to check to make sure i am right so far, and a little help!: rewrite y=sqrt of (25x-75)+3 to make it easier to graph and desribe the translation

OpenStudy (welshfella):

take 25 out of the parentheses:- = 5 sqrt(x -3) + 3

OpenStudy (anonymous):

i simplified it to where it is \[y=5\sqrt{(x-3)+3}\] but im confused how many units it translates and in which direction

OpenStudy (welshfella):

oh i see the square root covers the + 3 as well

OpenStudy (welshfella):

are you sure of that? - if so we can simplify the square root to sqrt x

OpenStudy (phi):

I hope the problem is \[ y = \sqrt{25x-75} + 3\] otherwise it's an ugly equation...

OpenStudy (phi):

if it is \[ y = \sqrt{25x-75} + 3 \] then as welsh posted, you have \[ y = 5 \sqrt{x-3}+3 \] the "parent graph" is \( y= \sqrt{x} \) the +3 shifts the graph up the (x-3) shifts the graph to the right 3 units the 5 makes the graph "steeper"

OpenStudy (anonymous):

yeah, it is. that was my bad

OpenStudy (anonymous):

and thank you to both of you !

OpenStudy (welshfella):

yw